OpenStudy (elleblythe):

How to solve sec165?

OpenStudy (mathmath333):

\(\large\bf\ sec(\theta+90)=-cosec(\theta)= \dfrac{-1}{sin(\theta)}\)

OpenStudy (cwrw238):

or if you want to use the unit circle